OpenStudy (anonymous):

A company manufacturers and sells x televisions per month. The monthly cost and and revenue equations are: c(x) = 82,000 +50x, for 0

OpenStudy (anonymous):

R(x)= - (1/30) x² + 500x R'(x) = - (1/15) x + 500 R'(x) = 0 --> x = 500 * 15 = 7,500 televisions

OpenStudy (anonymous):

-> Max Revenue: R ( 7, 500 ) = ....

OpenStudy (anonymous):

Would I plug in 7500 back to the original equation of 82,000 +50x, or is that something different?

OpenStudy (anonymous):

R ( x ) = - (1/30) x² + 500x R ( 7, 500 ) = - (1/30) 7, 500 ² + 500 *7, 500